# Can Large Language Models Improve Phishing Defense? A Large-Scale Controlled Experiment on Warning Dialogue Explanations

### v1

@@ 6.1 Internal Validity
Internal validity can be threatened by hidden factors that can compromise the achieved conclusions.
Confounding Variables. Latent factors such as prior knowledge and experience with phishing may have influenced behaviour and response to warnings. We used a between-subjects design to mitigate this threat, and participants were randomly assigned to conditions.
Deception and priming. At the beginning of the experiment, participants were informed that the study aimed to evaluate a new email client interface, thus hiding the true objective of the study, which was revealed at the end to avoid biasing participants towards unnatural behaviour. Even if this procedure was ethically approved to reduce demand characteristics, such framing may have influenced attentiveness and perceived task seriousness.
@@ 6.2 External Validity
Ecological validity. The email client we created mimics real and common features of web clients; however, users knew that they were involved in a study, thus in a safe environment. This reduced-risk setting can lead to unnatural behaviours that users exhibit when facing phishing attacks. This risk was mitigated by asking users to behave as naturally as possible.
Participant recruiting. To make the study population as representative as possible of a generic population to a phishing attack, we used Prolific by setting only gender and English knowledge as inclusion criteria. This guaranteed that participants were recruited worldwide and of different ages. However, considering the typical Prolific users, our sample may not include older adults, cybersecurity professionals, or individuals with low digital literacy.
@@ 6.3 Construct Validity
Behavioral measures. We adopted the click-through rate (CTR), the most widely adopted metric to measure the phishing vulnerability. However, clicks may occur in a controlled and safe setting for various reasons, including curiosity or misunderstanding. Thus, this behaviour may not always reflect a failure to detect phishing intent.
@@ 6.4 Statistical Validity
Multiple comparisons. In our study, different statistical tests such as Chi-square analyses, Kruskal-Wallis tests, and logistic regressions were conducted. However, while Bonferroni correction was applied where applicable, the risk of Type I errors was reduced, but it marginally remained.
@@ 7 Lessons learned
Triangulating qualitative and quantitative findings, we distilled lessons learned and directly provided guidance to design and implement future decisions. The four research questions are correlated with these lessons, thus determining the practical implications of our studies. They explain how the interaction between various design options, methods of explanation, and user differences affects the effectiveness and the perceived utility of AI-produced phishing warning messages.
Lesson 1 (RQ1): Automating warnings generation with LLMs without losing effectiveness
Feature-based explanations generated by LLM proved to be in line, or sometimes better, than the phishing protection of manually written explanations. To be more precise, the CTR of the Claude-generated explanations was even lower than that of the baseline and the Llama, thus indicating that a quality LLM-generated explanation is as effective as one written by experts in defending users.
These results align with past studies examining the automatic performance of explanations when used in security settings. For example, Chen et al. (2021) found users deemed automated phishing explanations beneficial and credible, but not all were protective. Similarly, Liao et al. (2020) showed that the overall meaning of an explanation, as opposed to its stylistic structure, has more influence on user trust and behaviour. All these results together demonstrate that properly fine-tuned LLM outputs exceed (or are on par with) human efforts in that area when considering causal reasoning.
Takeaway. If applying high-quality LLMs along with focused prompt engineering, explanation messages on phishing warnings can achieve the performance of an expert-created message, thus making it possible to create explanation systems at scale.
@@
In investigating feature-based explanation styles, the present study reveals that in the case of true positive emails, feature-based resulted in a lower CTR (11.03%) with respect to counterfactual ones (12.04%). This difference was more pronounced in the case of Claude-generated explanations (Claude FB: 7.48% vs. Claude CF: 10.27%), although this should be interpreted as an observed trend rather than a statistically significant effect, given the p-value exceeding conventional thresholds. These insights are consistent with the recent study by Zhao et al. (2024), showing that slight differences in language model tuning may trigger strong behavioural outcomes. The analysis of Item 9 further illustrates that explanations issued by Claude more frequently prompted interpretations invoking data theft or credential loss, thus suggesting a heightened salience of risk.
Takeaway. Feature-based explanations may help reduce clicks on phishing attempts, although further research is needed to deepen the investigation of this trend; in particular, LLMs such as Claude appear to convey a greater sense of threat.
@@ 7.1 Lesson 3 (RQ2): Counterfactuals may reduce false alarm overrides
Even if counterfactuals showed no overall CTR advantage, it emerged that they might reduce false positive overrides (10.00% vs. 14.06% for feature-based). Also, qualitative responses (e.g., “The email would have been safe if…” [P393]) indicate counterfactuals help users distinguish true threats from benign anomalies. Thus, counterfactuals help users reason about why the content may or may not be dangerous.
@@
The analysis of Item 9 revealed that feature-based and counterfactual explanations elicit different sources of threats in users. In the case of feature-based explanations, users tend to identify the outcome of the threat, for example, stolen credentials or leaked data. On the contrary, in the case of counterfactual explanations, users often pinpoint the mechanism of deception, especially deceptive links and their mismatch with the related labels. This finding aligns with findings in literature Djatsa (2020); Greco et al. (2025), which revealed that the explanations of the threat influence the users’ beliefs about where the risk resides.
@@
Lesson 5 (RQ3): Static metrics and user perceptions can signal warning efficacy.
Empirical evidence indicates that better readability scores (according to the SMOG index and Flesch-Kincaid grade level scores) are usually associated with low CTRs. Two contrasting sets of explanatory messages illustrate this phenomenon: Claude’s feature-based phrases, which achieved a SMOG score of 7.9 and a Flesch-Kincaid grade level of 5.6, registered a CTR of 9.39%, whereas Llama’s analogous constructs, with a SMOG score of 8.9 and a Flesch-Kincaid grade level of 6.9, demonstrated a higher CTR of 14.65%. The above observations provide tentative support for the hypothesis that adopting less complicated wording might enhance the effectiveness of the warning message.
Moreover, the questionnaire results showed that perceived trust and perceived risk were important predictors of CTR. In the false-positive subset, higher values of perceived trust in the warning significantly decreased the odds of clicking (OR=0.44, p=0.038), suggesting an overconfidence effect in the warning, which may lead to discarding genuine emails.
[…]
Takeaway. A combination of readability measures and user-based assessment, such as perceived trust, perceived risk, and message clarity, represents a promising approach to enhancing the efficacy of warnings and minimizing the vulnerability to phishing.
@@
The study results suggest that familiarity with warnings and the extended use of the internet daily are both highly significant predictors of clicking on phishing emails. In particular, the odds of clicking were more than doubled by a one-point increase in familiarity (OR = 2.20, p = .001), which indicates that habituation is a direct detriment to the effectiveness of warnings. Similarly, each extra hour spent online per day increased click probability by about 25% (OR = 1.25, p = .024), implying that individuals exposed to a lot of online content have a high chance of adopting automatic or shortcut browsing behaviours. These findings are consistent with prior studies on habituation in warning design literature (Amran et al., 2018, Akhawe and Felt, 2013, Egelman et al., 2008), which proved that familiarity may reduce people’s sensitivity to a familiar warning signal.
@@
The findings are congruent with past results by Sheahan et al. (2024) who state that constructive friction promotes reflection before doing something. Our finding is also in line with Kahneman (2011), which illustrates how shifting from fast (impulsive) to slow (deliberative) thinking improves judgment in risk-laden contexts.
@@
Logistic regressions revealed that gender had significant and systematic effects on click behaviour. The likelihood of female participants clicking on phishing emails was significantly lower, and the odds ratios averaged almost zero in true-positive and false-positive phishing email conditions. On the contrary, male participants with higher values of Need for Cognition (NFC) were over twice as likely to click in the false-positive conditions with counterfactual warnings (OR = 2.16, p = .012). The gender-by-NFC interaction indicates that some cognitive characteristics, in combination with gender, can increase vulnerability, perhaps due to over-trusting of the reasoning ability when subjected to ambiguous cues by the user. These results support the findings obtained by Greitzer et al. (2021), who also observed that male participants were more susceptible to phishing attacks in simulated attacks; they explained the trend with increased overconfidence or lowered caution.
@@ 8 Conclusions
This study offers empirical evidence that LLMs can automatically produce phishing warning explanations that are as effective as, and in certain circumstances, more effective than, those crafted by humans. Notably, warnings generated by Claude 3.5 Sonnet markedly reduced click-through rates on phishing links compared with Llama 3.3 70B and human-composed messages, especially when a feature-based explanation style was used. Counterfactual explanations showed initial potential in reducing false positives, yet feature-based explanations remained more reliable for identifying genuine threats.
The viability of LLM-generated warnings was supported with user-perception measures: participants found them clear, trustworthy, and actionable, especially those generated by Claude. The study also found that mental workload, gender, and familiarity with warning dialogues significantly affected the warning efficacy. These results stress the role of individualisation and flexibility of security interfaces.
Overall, employing LLMs in phishing-defence systems can improve the scalability, responsiveness, and user interaction without reducing effectiveness. Future work should examine real-time deployment, user-personalisation strategies, and broader applications of LLM-generated explanations in cybersecurity interfaces.
@@ Acknowledgments
CRediT authorship contribution statement
F. M. Cau: Conceptualization, Methodology, Formal analysis, Writing - Original Draft, Writing - Review & Editing, Visualization.
G. Desolda: Conceptualization, Methodology, Validation, Formal analysis, Investigation, Data Curation, Writing - Original Draft, Writing - Review & Editing, Supervision, Project administration, Funding acquisition.
F. Greco: Conceptualization, Methodology, Software, Formal analysis, Investigation, Resources, Data Curation, Writing - Original Draft, Writing - Review & Editing, Visualization.
L. D. Spano: Conceptualization, Methodology, Formal analysis, Writing - Original Draft, Writing - Review & Editing, Visualization, Supervision, Project administration, Funding acquisition.
L. Viganò: Conceptualization. Methodology, Formal analysis, Writing - Original Draft, Writing - Review & Editing, Supervision.
Appendix A Warning Messages

### v2

@@ 6.1 Internal Validity
Internal validity can be compromised by hidden factors that may threaten the conclusions drawn.
Confounding Variables. Latent factors, such as prior knowledge and experience with phishing, may have influenced behavior and responses to warnings. We employed a between-subjects design to mitigate this threat, and participants were randomly assigned to one of the conditions.
Deception and priming. At the beginning of the experiment, participants were informed that the study aimed to evaluate a new email client interface, thereby concealing the true objective of the study, which was revealed at the end to avoid biasing participants towards unnatural behavior. Even if this procedure was ethically approved to reduce demand characteristics, such framing may have influenced attentiveness and the perceived seriousness of the task.
@@ 6.2 External Validity
Ecological validity. The email client we created mimics real and common features of web clients; however, users were aware that they were participating in a study, thus in a safe environment. This reduced-risk setting can lead to unnatural behaviours that users exhibit when facing phishing attacks. This risk was mitigated by asking users to behave as naturally as possible.
Participant recruiting. To make the study population as representative as possible of a generic population affected by phishing attacks, we used Prolific by setting only gender and English proficiency as inclusion criteria. This guaranteed that participants were recruited worldwide and of different ages. However, considering the typical Prolific users, our sample may not include older adults, cybersecurity professionals, or individuals with low digital literacy.
@@ 6.3 Construct Validity
Behavioral measures. We adopted the click-through rate (CTR), the most widely adopted metric to measure phishing vulnerability. However, clicks may occur in a controlled and safe setting for various reasons, including curiosity or misunderstanding. Thus, this behaviour may not always reflect a failure to detect phishing intent.
@@ 6.4 Statistical Validity
Multiple comparisons. In our study, various statistical tests, including Chi-square analyses, Kruskal-Wallis tests, and logistic regressions, were conducted. However, while the Bonferroni correction was applied where applicable, the risk of Type I errors was reduced, but it marginally remained.
@@ 7 Lessons learned
Triangulating qualitative and quantitative findings, we distilled lessons learned and provided direct guidance on designing and implementing future decisions. The four research questions are correlated with these lessons, thus determining the practical implications of our studies. They explain how the interaction between various design options, methods of explanation, and user differences affects the effectiveness and the perceived utility of AI-produced phishing warning messages.
Lesson 1 (RQ1): The Economic Case for Automated Explanations
Our results provide empirical evidence that LLM-generated explanations—specifically those from Claude 3.5 Sonnet—achieve protection levels that are statistically indistinguishable from expert-crafted messages (p >> 0.05). Crucially, the observed click-through rate for Claude (9.39%) was numerically lower than the manual baseline (9.94%), indicating that automation introduces no performance degradation.
This finding carries substantial practical significance. Manually crafting context-aware explanations for every potential phishing vector is unscalable and cost-prohibitive. Conversely, our results demonstrate that LLMs can satisfy a requirement of operational non-inferiority purely through automated prompting. This shifts the paradigm from asking ”Are LLMs significantly better than experts?” to recognizing that ”LLMs are sufficiently capable to democratize high-quality security explanations at scale without compromising safety.” These findings align with recent work suggesting that properly fine-tuned LLM outputs can match human efforts in causal reasoning (Liao2020CHI), extending this validity to the critical domain of phishing defense.
Takeaway. Security practitioners can leverage high-quality LLMs to automate warning generation. Since the automated output proves to be statistically comparable to the manual ”gold standard,” the focus should shift from content creation to prompt engineering and validation, enabling scalable, real-time defenses against evolving threats.
@@
In investigating feature-based explanation styles, the present study reveals that in the case of true positive emails, feature-based resulted in a lower CTR (11.03%) with respect to counterfactual ones (12.04%). This difference was more pronounced in the case of Claude-generated explanations (Claude FB: 7.48% vs. Claude CF: 10.27%), although this should be interpreted as an observed trend rather than a statistically significant effect, given the p-value exceeding conventional thresholds. These insights are consistent with the recent study by zhao2024explainability, which shows that slight differences in language model tuning can trigger strong behavioral outcomes. The analysis of Item 9 further illustrates that explanations issued by Claude more frequently prompted interpretations invoking data theft or credential loss, thus suggesting a heightened salience of risk.
Takeaway. Feature-based explanations may help reduce clicks on phishing attempts, although further research is needed to investigate this trend more deeply; in particular, LLMs such as Claude appear to convey a greater sense of threat.
@@ 7.1 Lesson 3 (RQ2): Counterfactuals may reduce false alarm overrides
Even if counterfactuals showed no overall CTR advantage, it emerged that they might reduce false positive overrides (10.00% vs. 14.06% for feature-based). Additionally, qualitative responses (e.g., “The email would have been safe if…” [P393]) suggest that counterfactuals help users distinguish true threats from benign anomalies. Thus, counterfactuals help users reason about why the content may or may not be dangerous.
@@
The analysis of Item 9 revealed that feature-based and counterfactual explanations elicit different sources of threats in users. In the case of feature-based explanations, users tend to identify the outcome of the threat, for example, stolen credentials or leaked data. On the contrary, in the case of counterfactual explanations, users often pinpoint the mechanism of deception, particularly with deceptive links and their mismatch with related labels. This finding aligns with findings in the literature djatsa2020threat; greco2025enhancing, which reveal that the explanations of the threat influence users’ beliefs about where the risk resides.
@@
Lesson 5: The Readability-Security Paradox
Our findings reveal a counterintuitive divergence between perceived understandability and behavioral effectiveness, a phenomenon we term the Readability-Security Paradox. Llama 3.3 generated the most ”understandable” warnings (statistically significant vs. baseline, $p=0.002$) yet resulted in higher click-through rates. In contrast, Claude 3.5 produced more complex explanations—linking features to consequences, such as ”identity theft”—and achieved better protection despite lower readability scores. This suggests that clarity alone is insufficient for security warnings. Excessive simplification or a neutral tone (characteristic of Llama’s outputs) might inadvertently lower the user’s risk perception, leading to a ”compliance” failure despite ”comprehension” success. Effective security warnings may require a degree of ”cognitive friction” or semantic severity that purely readable text fails to convey. Future LLM prompting strategies should therefore prioritize ”risk salience” over simple ”readability”, tuning models to balance clarity with fear appeals or consequence-based framing.
Takeaway. Do not optimize solely for readability metrics. LLM prompts should be engineered to balance linguistic clarity with risk salience. Instruct models to explicitly articulate severe consequences (e.g., data theft) rather than maintaining a neutral tone, thereby introducing the necessary cognitive friction to drive compliant behavior.
@@
The study results suggest that familiarity with warnings and the extended use of the internet daily are both highly significant predictors of clicking on phishing emails. In particular, the odds of clicking were more than doubled by a one-point increase in familiarity (OR = 2.20, p = .001), which indicates that habituation is a direct detriment to the effectiveness of warnings. Similarly, each extra hour spent online per day increased click probability by about 25% (OR = 1.25, p = .024), implying that individuals exposed to a lot of online content have a high chance of adopting automatic or shortcut browsing behaviours. These findings are consistent with prior studies on habituation in the warning design literature (amran2018habituation, akhawe2013alice, egelman2008warned), which have shown that familiarity may reduce people’s sensitivity to a familiar warning signal.
@@
The findings are congruent with past results by (sheahan2024designing) who state that constructive friction promotes reflection before doing something. Our finding is also in line with (Kahneman2011Thinking), which illustrates how shifting from fast (impulsive) to slow (deliberative) thinking improves judgment in risk-laden contexts.
@@
Logistic regressions revealed that gender had a significant and systematic effect on click behavior. The likelihood of female participants clicking on phishing emails was significantly lower, and the odds ratios averaged almost zero in true-positive and false-positive phishing email conditions. On the contrary, male participants with higher values of Need for Cognition (NFC) were over twice as likely to click in the false-positive conditions with counterfactual warnings (OR = 2.16, p = .012). The gender-by-NFC interaction suggests that certain cognitive characteristics, in combination with gender, may increase vulnerability, possibly due to an over-trust in the reasoning ability when exposed to ambiguous cues by the user. These results support the findings obtained by greitzer2021experimental, who also observed that male participants were more susceptible to phishing attacks in simulated attacks; they explained the trend with increased overconfidence or lowered caution.
@@
Lesson 9: Safety Guardrails and the Risk of Unconstrained Generation
The selection phase of the LLM models revealed a critical insight: unconstrained LLMs, i.e., models generating text without external verification layers, pose significant security risks. In our preliminary tests, models such as Gemini 1.5 Flash and Llama 3.1 8B hallucinated when prompted directly, generating nonexistent URL features or providing vague safety assurances for malicious links. This result indicates that purely generative approaches are insufficient for critical defense tasks. To ensure safety and mitigate this risk, our study employed a Human-in-the-Loop validation approach. Similarly, the real-world deployment of the final system must consider a ”hybrid architecture” where generative outputs are constrained by deterministic checks (e.g., code that verifies whether the explained URL actually matches the email header) before being presented to the user.
Takeaway. Generative AI for warning dialogs requires a ”trust but verify” approach. Outputs generated by LLMs must be rigorously validated to prevent hallucinations, ensuring that false assurances do not paradoxically lower user protection.
@@ 8 Conclusions
This study provides empirical evidence that LLMs can automatically generate phishing warning explanations that are as effective as, and in certain cases, more effective than those crafted by humans. Notably, warnings generated by Claude 3.5 Sonnet markedly reduced click-through rates on phishing links compared with Llama 3.3 70B and human-composed messages, especially when a feature-based explanation style was used. Counterfactual explanations initially showed potential in reducing false positives; however, feature-based explanations remained more reliable for identifying genuine threats.
The viability of LLM-generated warnings was supported with user-perception measures: participants found them clear, trustworthy, and actionable, especially those generated by Claude. The study also found that mental workload, gender, and familiarity with warning dialogues significantly affected the warning efficacy. These results emphasize the importance of individualization and flexibility in security interfaces.
Overall, employing LLMs in phishing defense systems can improve scalability, responsiveness, and user interaction without reducing effectiveness. Future work should examine real-time deployment, user-personalisation strategies, and broader applications of LLM-generated explanations in cybersecurity interfaces.
@@
Data Availability Statement
All materials supporting the findings of this study are publicly available in the following repository:
The repository includes: (i) the complete source code of the experimental web platform used to conduct the user study; (ii) detailed documentation of the explanation-generation process, including all feature-based and counterfactual explanations evaluated prior to the user study; (iii) screenshots of all email stimuli presented to participants during the experiment; and (iv) raw, anonymized interaction data together with processed datasets and scripts used for the statistical analyses reported in this paper.
All user data have been fully anonymized, and no personally identifiable information is included. The shared materials enable full transparency and support the reproducibility of the experimental design, data analysis, and reported results.
@@ Acknowledgments
{appendices}
9 A. Prompts for generating explanations
To generate the explanations used in the study, we used a modified version of APOLLO desolda2025apollo, an LLM-based tool for phishing email classification and explanation. While the full details of APOLLO are reported in the original paper desolda2025apollo, for the sake of auto-consistency, we provide the prompts that the system uses in the following.
9.1 Email classification prompt
Listing 1: First prompt. This makes the LLM generate a classification outcome for an email and generate a first, raw explanation.
You are a cybersecurity and human-computer interaction expert who has the goal of detecting if an email is legitimate or phishing and helping the user understand why a specific email is dangerous (or genuine) in order to make more informed decisions.
The user will submit the email (headers + subject + body) optionally accompanied by information on the URLs in the email as:
- server location;
- VirusTotal scans reporting the number of scanners that detected the URL as harmless, undetected, malicious.
Your goal is to output a JSON object containing:
- The classification result (label).
- The probability as a percentage of the email being phishing (0%=email is surely legitimate, 100%=email is surely phishing) (phishing_probability).
- A list of persuasion principles that were applied by the alleged attacker (if any); each persuasion principle should be an object containing:
the persuasion principle name (authority, scarcity, etc.),
the part of the email that makes you say that the persuasion principle is being applied;
a brief rationale for each principle.
- A list of 3 to 5 features that could indicate the danger (or legitimacy) of the email; the explanations must be understandable by users with no cybersecurity or computer expertise.
Desired format:
label: <phishing/legit>
phishing_probability: <0-100%>
persuasion_principles: [array of persuasion principles, each having: (name, specific sentences, rationale)]
explanation: [array of 3-5 features explained]
Email:
[…]
[BODY]
{email body}
[…]
URL Information:
Server location: {URL geolocation}
VirusTotal scan: [
harmless: {n_harmless},
undetected: {n_undetected},
malicious: {n_malicious}
9.2 Explanation generation prompts
The results of the email classification are then sent to the LLM, along with a second prompt, adapted from APOLLO desolda2025apollo, which varies according to the explanation type (feature-based or counterfactual).
Listing 2: This prompt generates a refined feature-based explanation message following a specific structure. Note that, for the aims of this study, the model is primed to explain a specific feature injected in the prompt (”feature to explain”).
Now consider the [feature to explain] feature and construct a brief explanation message construct a brief explanation message (max 50 words) directed to naive users (with no knowledge of cybersecurity) that will follow this structure:
1. description of the most relevant phishing feature
2. explanation of the hazard
3. consequences of a successful phishing attack
For example, an explanation that explains that the domain of a website is suspiciously young would be:
"The URL in the email leads to a website created N days ago. Young websites are famous for criminal activity. There is a potential risk if you proceed."
Another example of explaining that the email is suspicious because of too many special characters in its body would be:
"Many special characters have been detected in the email. Malicious people use them to disguise text and deceive you. Your data could be stolen."
Desired format:
[description of the feature]. [hazard explanation]. [consequences of a successful attack].
Listing 3: This prompt generates a refined counterfactual explanation message following a specific structure. Note that, for the aims of this study, the model is primed to explain a specific feature injected in the prompt (”feature to explain”).
Now consider the [feature to explain] feature and construct a brief counterfactual explanation message (max 50 words) directed to naive users (with no knowledge of cybersecurity or computers) that will follow this structure:
1. description of the phishing feature
2. hypothetical conditions for which it would have been safe
3. counter-example of a safe feature
Avoid using technical jargon such as "IP address", "domain", "URL", etc. It might also be useful to display the URL within the warning message.
For example, an explanation that explains that the domain of a website is suspiciously young would be:
"The URL in the email leads to a website created N days ago. The email would have been considered safe if the website were older than a typical attack cycle, as this timeframe provides a verifiable reputation. A safe link might point to a domain with years of established activity."
Another example of explaining that the email is suspicious because of too many special characters in its body would be:
"Many special characters have been detected in the email. The email would have been considered safe if it mostly contained standard text (letters and numbers), as this eliminates the possibility of visually deceptive text used to trick users. A safe email would only use characters typically found in formal correspondence."
Desired format:
[description of the feature]. [hypothetical safety conditions]. [counter-example]
10 B. Warning Messages